# Universal properties of the evolution of the Universe in modified loop quantum cosmology

### v2

@@ III.3.2 Polynomial of the first kind
where $V_{0}$ and $\mu$ are two constants. In this paper, we take the values of $V_{0}$ and $\mu$ as $V_{0}=3.3787\times 10^{-15}$ GeV and $\mu=0.31075$ Bhattacharya et al. (2023). Then, in Figs. 13 and 14 we plot the e-fold $N$ and the equation of state $w(t)$ for both $\dot{\phi}_{B}>0$ and $\dot{\phi}_{B}<0$.
Similar to the previous cases, the evolution of the universe in each of the pre- and post-bounce regimes can be universally divided into three epochs, describing by Eqs.(II) and (III.1), respectively.
In Fig. 15, we plot the functions $a(t)$ and $\phi(t)$ with various initial values of $\phi_{B}$ for $\dot{\phi}_{B}>0$ and $\dot{\phi}_{B}<0$, respectively.
@@ III.3.3 Polynomial of the second kind
In Figs. 16 and 17 we plot $N(\phi_{B})$ and the corresponding equation of state $w(t)$ for several initial conditions $\phi_{B}$ for $\dot{\phi}_{B}>0$ and $\dot{\phi}_{B}<0$, respectively. These values of $\phi_{B}$ guarantee that the e-folds during inflation will be about 60 as shown in Fig. 16. On the other hand, Fig. 17 shows again that the evolution of the universe in each of the pre- and post-bounce regimes can be universally divided into three different epochs.
@@ III.3.4 Generalized T-models
In Figs. 19 and 20 we plot $N(\phi_{B})$ and the corresponding equation of state $w(t)$ for several initial conditions $\phi_{B}$ with $\dot{\phi}_{B}>0$ and $\dot{\phi}_{B}<0$, respectively. These values of $\phi_{B}$ also guarantee that the e-folds of the expansions of the universe during the inflationary period will be about 60, as shown in Fig. 20. On the other hand, Fig. 20 shows again that the evolution of the universe in each of the pre- and post-bounce regimes can be universally divided into three different epochs, as described by Eq.(II) in the post-bounce regime and by Eq.(III.1) in the pre-bounce regime.
In Fig. 21, we plot the numerical solutions of $a(t)$ and $\phi(t)$ with several initial conditions for both $\dot{\phi}_{B}>0$ and $\dot{\phi}_{B}<0$ case, which all satisfy the condition (1.1) at the bounce. Again, they show the universality of the solutions.
@@ III.4 Natural Inflation
where $\mu\equiv f/{\cal{N}}$, $\Lambda\simeq m_{\text{GUT}}\simeq 10^{16}$ GeV, $f\simeq m_{\text{pl}}$, and ${\cal{N}}$ is a dimensionless constant. In the following, we consider only three representative cases, $\mu/m_{\text{pl}}=\left(0.3,\;1,\;5\right)$.
III.4.1 $\mu=0.3\;m_{\text{pl}}$
In this case, we first identify the values of $\phi_{B}$ with both $\dot{\phi}_{B}>0$ and $\dot{\phi}_{B}<0$, for which the e-folds can be approximately 60, as shown in Fig. 22.
@@
Figure 22: E-folds of natural inflation with the potential given by Eq.(3.15) and $\Lambda=10^{16}$ GeV and $\mu=0.3m_{\text{pl}}$ for different values of the initial conditions $\phi_{B}$, where (a) for the case $\dot{\phi}_{B}>0$ and (b) for the case $\dot{\phi}_{B}<0$. Left Panel: $\dot{\phi}_{B}>0$. Right Panel: $\dot{\phi}_{B}<0$.
@@
Figure 23: The equation of state $w(t)$ with different values of $\phi_{B}$ for for the natural inflation with the potential given by Eq.(3.15) and $\Lambda=10^{16}$ GeV and $\mu=0.3m_{\text{pl}}$, where (a) $\dot{\phi}_{B}>0$ and (b) $\dot{\phi}_{B}<0$.
@@
Refer to caption Refer to caption
Figure 24: The numerical solutions of $a(t)$ and $\phi(t)$ for various initial conditions of $\phi_{B}$ for the natural inflation with the potential given by Eq.(3.15) and $\Lambda=10^{16}$ GeV and $\mu=0.3$, where (a) for $\dot{\phi}_{B}>0$, and (b) for $\dot{\phi}_{B}>0$.
On the other hand, in Fig. 23 the equation of state is plotted with some of these initial values (with which the corresponding e-folds are about 60). From the figure, it can be seen again that the evolution of the universe is universal in both the pre- and post-bounce regimes, and in each of which three distinguishable epochs exist, as described by Eqs.(II) and (III.1).
The corresponding numerical solutions of $a(t)$ and $\phi(t)$ are shown in Fig. 24 for $\dot{\phi}_{B}>0$ and $\dot{\phi}_{B}<0$, respectively, form which it cna be seen clearly the universality of the numerical solutions.
Refer to caption Refer to caption
Figure 25: E-folds of natural inflation with the potential given by Eq.(3.15) and $\Lambda=10^{16}$ GeV and $\mu=m_{\text{pl}}$ for different values of the initial conditions $\phi_{B}$. Left Panel: $\dot{\phi}_{B}>0$. Right Panel: $\dot{\phi}_{B}<0$.
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 26: The equation of state $w(t)$ with different values of $\phi_{B}$ for for the natural inflation with the potential given by Eq.(3.15) and $\Lambda=10^{16}$ GeV and $\mu=1\;m_{\text{pl}}$, where (a) $\dot{\phi}_{B}>0$ and (b) $\dot{\phi}_{B}<0$.
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 27: The numerical solutions of $a(t)$ and $\phi(t)$ for various initial conditions of $\phi_{B}$ for the natural inflation with the potential given by Eq.(3.15) and $\Lambda=10^{16}$ GeV and $\mu=1m_{\text{pl}}$, where (a) for $\dot{\phi}_{B}>0$ and (a) for $\dot{\phi}_{B}<0$.
III.4.2 $\mu=m_{\text{pl}}$
In this subsection, similar to the previous ones, we first identify the initial conditions $\phi_{B}$ for both $\dot{\phi}_{B}>0$ and $\dot{\phi}_{B}<0$, with which the e-folds of inflation are approximately 60. In Fig. 25 we show two sets of such data. Note that the number of such sets is infinitely large because the potential is now a periodic function of $\phi$. With some of the initial conditions presented in Fig. 25, we plot the corresponding equation of state $w(t)$ in Fig. 26, which again shows the universal properties described by Eqs.(II) and (III.1).
Conversely, the corresponding numerical solutions of $a(t)$ and $\phi(t)$ are given in Fig. 27 for $\dot{\phi}_{B}>0$ and $\dot{\phi}_{B}<0$, respectively.
Refer to caption Refer to caption
Figure 28: E-folds of natural inflation with the potential given by Eq.(3.15) and $\Lambda=10^{16}$ GeV and $\mu=5m_{\text{pl}}$ for different values of the initial conditions $\phi_{B}$. Left Panel: $\dot{\phi}_{B}>0$. Right Panel: $\dot{\phi}_{B}<0$.
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 29: The equation of state $w(t)$ with different values of $\phi_{B}$ for for the natural inflation with the potential given by Eq.(3.15) and $\Lambda=10^{16}$ GeV and $\mu=5\;m_{\text{pl}}$, where (a) $\dot{\phi}_{B}>0$ and (b) $\dot{\phi}_{B}<0$.
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 30: The numerical solutions of $a(t)$ and $\phi(t)$ for various initial conditions of $\phi_{B}$ with $\dot{\phi}_{B}>0$ for natural inflation with the potential given by Eq.(3.15) and $\Lambda=10^{16}$ GeV and $\mu=5m_{\text{pl}}$.
III.4.3 $\mu=5\;m_{\text{pl}}$
In this case, in Fig. 28 we plot the e-folds of the natural inflation for both $\dot{\phi}_{B}>0$ and $\dot{\phi}_{B}<0$, whereby we can read off the initial conditions of $\phi_{B}$ for which the inflation will last about 60 e-folds.
Then, with several choices of such initial values of $\phi_{B}$ presented in Fig. 28, the equation of state $w(t)$ is plotted in Fig. 29. Again, we can clearly identify the three distinguishable epochs in each of the pre- and post-bounce regimes.
The corresponding numerical solutions of $a(t)$ and $\phi(t)$ are plotted in Fig. 30 for both $\dot{\phi}_{B}>0$ and $\dot{\phi}_{B}<0$.
@@ IV Analytical Solutions for the evolution of the Universe
$\phi(t)=e_{3}+\text{sgn}\left(\dot{\phi}_{B}\right)(e_{0}+e_{2}t)e^{e_{1}t},\;(t\leq t_{B}),$ (3.19)
where $e_{m}\;(m=0,...,3)$ are the fitting constants. Again, we require that $\phi(t)$ and its first derivative be continuous across the bounce, where $\phi(t)$ is given by Eq.(II) in the post-bounce regime ($t\geq t_{B}$). Then, we find that with the choices of fitting constants as
$e_{0}$ $\simeq$ $0.353,\;\;\;e_{1}\simeq 0.981,$
$e_{2}$ $\simeq$ $0.943,\;\;\;e_{3}\simeq-1.704,$ (3.20)
@@
It should be noted that
$e_{3}=\phi_{\infty}-\phi_{B},$ (3.22)
$e_{0}=\text{sgn}\left(\dot{\phi}_{B}\right)\left(2\phi_{B}-\phi_{\infty}\right),$
where $\phi_{\infty}\equiv\phi(t=-\infty)$. Therefore, $e_{0}$ and $e_{3}$ depends on the choice of the initial values of $\phi_{B}$ as well as the sign of $\dot{\phi}_{B}$, as can be seen from Eqs.(IV) and (IV). However, once the sign of $\dot{\phi}_{B}$ is chosen, the coefficient $e_{3}$ is independent of the inflationary potentials, as can be seen from Figs. 31 and 32.
@@ V Conclusions and Remarks
Moreover, the numerical solutions with the initial conditions that satisfy Eq.(1.1) considered in all the models of this paper are universal, as shown explicitly by Fig. 31, and can be approximated well by analytical solutions cast in the forms (3.16) for the expansion factor $a(t)$ and (3.19) for the inflation $\phi(t)$, where the constants $d_{n}$ and $e_{n}$ appearing in these expressions cna be obtained by fititng the analytical solutions with the numerical ones. For the expansion factor $a(t)$ we have found that the fitting ocnstants $d_{n}$ are given by Eq.(IV) for both $\dot{\phi}_{B}>0$ and $\dot{\phi}_{B}<0$. However, for the scalar field, the fitting constant $e_{n}$ depend on the signs of $\dot{\phi}_{B}$, and are given respectively by Eqs.(IV) and (IV). This is because these coefficients depend on the choice of the initial values of $\phi_{B}$ and the signs of $\dot{\phi}_{B}$, as shown by Eq.(3.22).
The above remarkable features are also shared in the evolution of the Universe in the post-bounce regime not only in the framework of mLQC-I Li et al. (2018a, b, 2019a, 2021) but also in that of LQC Zhu et al. (2017); Shahalam et al. (2017, 2018); Sharma et al. (2018, 2019); Shahalam et al. (2020). We believe that this universal behavior of background evolution will significantly simplify the studies of cosmological perturbations in mLQC-I, as has already been the case in LQC Zhu et al. (2017).

### v3

@@ III.3.2 Polynomial of the first kind
where $V_{0}$ and $\mu$ are two constants. In this paper, we take the values of $V_{0}$ and $\mu$ as $V_{0}=3.3787\times 10^{-15}$ GeV and $\mu=0.31075$ Bhattacharya et al. (2023).
Note that, similar to the chaotic potential, now the polynomial potential of the first kind has the symmetry
$V(\phi)=V(-\phi).$ (3.13)
As a results, the case with $\dot{\phi}_{B}<0$ can be obtained from the case with $\dot{\phi}_{B}>0$ via the symmetry (3.7). Clearly this is true for all the cases with the above symmetry, including the polynomial potential of the second kind and the generalized T-models to be considered below. Therefore, in all these cases, we shall consider only the cases with $\dot{\phi}_{B}>0$ without furthering noticing it.
Then, in Figs. 13 and 14 we plot the e-fold $N$ and the equation of state $w(t)$ for $\dot{\phi}_{B}>0$. Similar to the previous cases, the evolution of the universe in each of the pre- and post-bounce regimes can be universally divided into three epochs, describing by Eqs.(II) and (III.1), respectively. In Fig. 15, we plot the functions $a(t)$ and $\phi(t)$ with various initial values of $\phi_{B}$ for $\dot{\phi}_{B}>0$.
@@ III.3.3 Polynomial of the second kind
In Figs. 16 and 17 we plot $N(\phi_{B})$ and the corresponding equation of state $w(t)$ for several initial conditions $\phi_{B}$ for $\dot{\phi}_{B}>0$. These values of $\phi_{B}$ guarantee that the e-folds during inflation will be about 60 as shown in Fig. 16. On the other hand, Fig. 17 shows again that the evolution of the universe in each of the pre- and post-bounce regimes can be universally divided into three different epochs.
@@ III.3.4 Generalized T-models
In Figs. 19 and 20 we plot $N(\phi_{B})$ and the corresponding equation of state $w(t)$ for several initial conditions $\phi_{B}$ with $\dot{\phi}_{B}>0$. These values of $\phi_{B}$ also guarantee that the e-folds of the expansions of the universe during the inflationary period will be about 60, as shown in Fig. 20. On the other hand, Fig. 20 shows again that the evolution of the universe in each of the pre- and post-bounce regimes can be universally divided into three different epochs, as described by Eq.(II) in the post-bounce regime and by Eq.(III.1) in the pre-bounce regime.
In Fig. 21, we plot the numerical solutions of $a(t)$ and $\phi(t)$ with several initial conditions for $\dot{\phi}_{B}>0$, which all satisfy the condition (1.1) at the bounce. Again, they show the universality of the solutions.
@@ III.4 Natural Inflation
where $\mu\equiv f/{\cal{N}}$, $\Lambda\simeq m_{\text{GUT}}\simeq 10^{16}$ GeV, $f\simeq m_{\text{pl}}$, and ${\cal{N}}$ is a dimensionless constant. Because of the periodic property of the potential, without loss of the generality, we can restrict ourselves only to the region
$-\pi\leq\frac{\phi}{\mu}\leq\pi.$ (3.17)
In addition, the potential is also symmetric under the replacement $\phi\rightarrow-\phi$. So, in this case it is also sufficient to consider only the case $\dot{\phi}_{B}>0$. In the following, we consider only two representative cases, $\mu/m_{\text{pl}}=\left(1,\;5\right)$.
Figure 22: E-folds of natural inflation with the potential given by Eq.(3.16), $\dot{\phi}_{B}>0$, $\Lambda=10^{16}$ GeV, and $\mu=m_{\text{pl}}$ for different values of the initial conditions $\phi_{B}$.
@@
Figure 23: The equation of state $w(t)$ with different values of $\phi_{B}$ for for the natural inflation with the potential given by Eq.(3.16), $\dot{\phi}_{B}>0$, $\Lambda=10^{16}$ GeV and $\mu=1\;m_{\text{pl}}$.
@@
Figure 24: The numerical solutions of $a(t)$ and $\phi(t)$ for various initial conditions of $\phi_{B}$ for the natural inflation with the potential given by Eq.(3.16), $\dot{\phi}_{B}>0$, $\Lambda=10^{16}$ GeV and $\mu=1m_{\text{pl}}$.
III.4.1 $\mu=m_{\text{pl}}$
In this subsection, we first identify the initial conditions $\phi_{B}$ for both $\dot{\phi}_{B}>0$ and $\dot{\phi}_{B}<0$, with which the e-folds of inflation are approximately 60. In Fig. 22 we show two sets of such data. Note that the number of such sets is infinitely large because the potential is now a periodic function of $\phi$. With some of the initial conditions presented in Fig. 22, we plot the corresponding equation of state $w(t)$ in Fig. 23, which again shows the universal properties described by Eqs.(II) and (III.1). The corresponding numerical solutions of $a(t)$ and $\phi(t)$ are given in Fig. 24 for $\dot{\phi}_{B}>0$ and $\dot{\phi}_{B}<0$, respectively.
Figure 25: E-folds of natural inflation with the potential given by Eq.(3.16), $\dot{\phi}_{B}>0$, $\Lambda=10^{16}$ GeV and $\mu=5m_{\text{pl}}$ for different values of the initial conditions $\phi_{B}$.
@@ III.4.1 $\mu=m_{\text{pl}}$
Figure 26: The equation of state $w(t)$ with different values of $\phi_{B}$ for for the natural inflation with the potential given by Eq.(3.16), $\dot{\phi}_{B}>0$, $\Lambda=10^{16}$ GeV and $\mu=5\;m_{\text{pl}}$.
@@
Figure 27: The numerical solutions of $a(t)$ and $\phi(t)$ for various initial conditions of $\phi_{B}$ with $\dot{\phi}_{B}>0$ for natural inflation with the potential given by Eq.(3.16), $\dot{\phi}_{B}>0$, $\Lambda=10^{16}$ GeV and $\mu=5m_{\text{pl}}$.
III.4.2 $\mu=5\;m_{\text{pl}}$
In this case, in Fig. 25 we plot the e-folds of the natural inflation for both $\dot{\phi}_{B}>0$ and $\dot{\phi}_{B}<0$, whereby we can read off the initial conditions of $\phi_{B}$ for which the inflation will last about 60 e-folds. Then, with several choices of such initial values of $\phi_{B}$ presented in Fig. 25, the equation of state $w(t)$ is plotted in Fig. 26. Again, we can clearly identify the three distinguishable epochs in each of the pre- and post-bounce regimes. The corresponding numerical solutions of $a(t)$ and $\phi(t)$ are plotted in Fig. 27 for both $\dot{\phi}_{B}>0$ and $\dot{\phi}_{B}<0$.
@@ IV Analytical Solutions for the evolution of the Universe
$\phi(t)-\phi_{B}=e_{3}+\text{sgn}\left(\dot{\phi}_{B}\right)(e_{0}+e_{2}t)e^{e_{1}t},\;(t\leq t_{B}),$ (3.21)
where $e_{m}\;(m=0,...,3)$ are the fitting constants. Again, we require that $\phi(t)$ and its first derivative be continuous across the bounce, where $\phi(t)$ is given by Eq.(II) in the post-bounce regime ($t\geq t_{B}$). Then, fitting the above with the numerical data, we find that with the choices of fitting constants as
$e_{0}$ $\simeq$ $0.3526,\;\;\;e_{1}\simeq 0.9808,$
$e_{2}$ $\simeq$ $0.0943,\;\;\;e_{3}\simeq-0.3526,$ (3.22)
@@
To understanding the above fitting results, let us first note that at the bounce $t=0$ and at the remote contracting phase $t\rightarrow-\infty$, we have
$0=e_{3}+\text{sgn}\left(\dot{\phi}_{B}\right)e_{0},$
$\phi_{-\infty}-\phi_{B}=e_{3},$ (3.24)
where $\phi_{-\infty}\equiv\phi(t=-\infty)$. On the other hand, from Figs. 27 and 28, we also have $\phi_{-\infty}-\phi_{B}=-\text{sgn}\left(\dot{\phi}_{B}\right){\cal{A}}_{0}$, where ${\cal{A}}_{0}\simeq 0.35326$. Thus, finally we have
$e_{0}={\cal{A}}_{0}\simeq 0.3526,\quad e_{3}=-\text{sgn}\left(\dot{\phi}_{B}\right){\cal{A}}_{0}.$ (3.25)
@@ V Conclusions and Remarks
Moreover, the numerical solutions with the initial conditions that satisfy Eq.(1.1) considered in all the models of this paper are universal, as shown explicitly by Fig. 28, and can be approximated well by analytical solutions cast in the forms (3.18) for the expansion factor $a(t)$ and (3.21) for the inflation $\phi(t)$, where the constants $d_{n}$ and $e_{n}$ appearing in these expressions cna be obtained by fitting the analytical solutions with the numerical ones. For the expansion factor $a(t)$ we have found that the fitting constants $d_{n}$ are given by Eq.(IV) for both $\dot{\phi}_{B}>0$ and $\dot{\phi}_{B}<0$. However, for the scalar field, the fitting constant $e_{n}$ depend on the signs of $\dot{\phi}_{B}$, and are given respectively by Eqs.(IV) and (IV).
The above remarkable features are also shared in the evolution of the Universe in the post-bounce regime not only in the framework of mLQC-I Li et al. (2018a, b, 2019a, 2021) but also in that of LQC Zhu et al. (2017a, b); Shahalam et al. (2017, 2018); Sharma et al. (2018, 2019); Shahalam et al. (2020). We believe that this universal behavior of background evolution will significantly simplify the studies of cosmological perturbations in mLQC-I, as has already been the case in LQC Zhu et al. (2017b).